# The (R)Evolution of Multimodal Large Language Models: A Survey

### v1

@@ 2.3 Vision-to-Language Adapters
Q-Former. It is a Transformer-based model proposed in BLIP-2 Li et al. (2023f) and then used in several other approaches Chen et al. (2023d); Dai et al. (2023); Hu et al. (2024). It is characterized by its adaptable architecture, which consists of two Transformer blocks that share mutual self-attention layers, facilitating the alignment process between visual and textual representations. It involves a set of learnable queries that interact within the self-attention layers and interface with visual features via a cross-attention mechanism. Textual and visual elements communicate via shared self-attention within the modules.
@@
Additional Cross-Attention Layers. This approach has been proposed in Flamingo Alayrac et al. (2022) with the integration of dense cross-attention blocks among the existing pre-trained layers of the LLM. The newly added layers are often combined with a zero-initialized tanh-gating mechanism to ensure that, upon initialization, the conditioned model acts as its original version. The use of additional cross-attention layers imposes the need to train them from scratch, increasing the number of trainable parameters compared to other alternatives. To reduce computational complexity, this strategy is usually paired with a Perceiver-based component Jaegle et al. (2021) that reduces the number of visual tokens before they are fed to the LLM. Since its introduction, several models Awadalla et al. (2023); Chen et al. (2023b); Laurençon et al. (2023); Li et al. (2023a) employ this technique to connect the visual modality with the underlying LLM, demonstrating enhanced training stability and improved performance.
@@ 2.4 Multimodal Training
Single-Stage Training. This possibility is explored by LLaMA-Adapter Gao et al. (2023) which introduces additional trainable parameters to encapsulate the visual knowledge and manage text-only instruction learning at the same time. To achieve this, the model undergoes joint training using image-text pairs and instructions, operating on separate parameters. Concurrently, the model proposed in Koh et al. (2023b) adapts the final loss function by incorporating two contrastive losses for image-text retrieval. During the training, only three linear layers are updated. On a different line, Kosmos-1 Huang et al. (2023a) considers a frozen visual backbone and trains the language model of 1.3B parameters from scratch.
Flamingo Alayrac et al. (2022) and its open source variants Awadalla et al. (2023); Laurençon et al. (2023), instead, train the cross-attention layers and the Perceiver-based component to connect the visual features with the frozen LLM blocks. Additionally, Otter Li et al. (2023a) extends Flamingo’s training to increment its in-context capabilities.
Given the amount of training data currently available, approaches like SPHINX-X Gao et al. (2024) opt to perform a single all-in-one training stage in which to update all model components, possibly also using text-only data to preserve the conversational capabilities of the LLM.
@@
Training Data. During the first (or single) training stage, image-text pairs from different sources are usually employed, using datasets like LAION-2B Schuhmann et al. (2022), LAION-400M Schuhmann et al. (2021), Conceptual Captions Sharma et al. (2018), COYO-700M Byeon et al. (2022), and DataComp Gadre et al. (2023). Some methods Lin et al. (2023a) use these in combination with one or more datasets characterized by text interleaved with images usually scraped from the web, such as WebLI Chen et al. (2023i), MMC4 Zhu et al. (2023d), MMDialog Feng et al. (2023b), and OBELICS Laurençon et al. (2023). To address bias and noise in the previous datasets, StableLLaVA Li et al. (2023h) introduces newly collected data to be used during the first stage. This approach utilizes ChatGPT to generate data comprising image-generating prompts and content-based dialogues, and Stable Diffusion Rombach et al. (2022) to generate the corresponding images.
The subsequent stage, instead, leverages datasets for visual instruction tuning. Among them, the commonly used LLaVA-Instruct Liu et al. (2023e) extends COCO Lin et al. (2014) with GPT-4 generated instructions. Following this trend, Zhao et al. (2023a) scale-up dimensions by incorporating both manually and generated data, with high quality and diversity. Moreover, other multi-turn dialogue datasets have been proposed such as the one introduced in Dai et al. (2023) that converts 26 publicly available datasets into their visual instruction-following versions, LRV-Instruction Liu et al. (2023c) that aims to reduce hallucinations with more robust instructions, and LLaVAR Zhang et al. (2023h) that instead focuses on text-rich images.
@@
ContextDET Zang et al. (2023) ${}^{\bigstar}$ Swin-B - Visual Dialogue, VQA, Captioning, Detection, REC, RES
DetGPT Pi et al. (2023) ${}^{\bigstar}$ EVA ViT-g ${}^{\bigstar}$ Visual Dialogue, Detection
VisionLLM Wang et al. (2023e) ${}^{\blacktriangle}$ Intern-H ${}^{\blacktriangle}$ VQA, Captioning, Detection, Segmentation, REC
BuboGPT Zhao et al. (2023c) ${}^{\bigstar}$ EVA ViT-g RAM, G-DINO, ${}^{\bigstar}$ Visual Dialogue, Audio Understanding, Captioning, GroundCap
ChatSpot Zhao et al. (2023b) ${}^{\blacklozenge}$ CLIP ViT-L - Visual Dialogue, VQA, Captioning, Referring
GPT4RoI Zhang et al. (2023f) ${}^{\blacklozenge}$ OpenCLIP ViT-H - Visual Dialogue, VQA, Captioning, Referring
ASM Wang et al. (2023d) ${}^{\blacktriangle}$ EVA ViT-g - VQA, Captioning, Referring
LISA Lai et al. (2023) ${}^{\blacktriangle}$ CLIP ViT-L ${}^{\blacklozenge}$ Visual Dialogue, Captioning, RES
PVIT Chen et al. (2023a) ${}^{\blacklozenge}$ CLIP ViT-L ${}^{\bigstar}$ Visual Dialogue, VQA, Captioning, Referring
GLaMM Rasheed et al. (2023) ${}^{\blacktriangle}$ OpenCLIP ViT-H ${}^{\blacklozenge}$ Visual Dialogue, Captioning, Referring, REC, RES, GroundCap
Griffon Zhan et al. (2023) ${}^{\blacklozenge}$ CLIP ViT-L - REC, Detection, Phrase Grounding
LLaFS Zhu et al. (2023c) ${}^{\blacktriangle}$ CLIP RN50 - Few-Shot Segmentation
NExT-Chat Zhang et al. (2023a) ${}^{\blacklozenge}$ CLIP ViT-L ${}^{\blacklozenge}$ Visual Dialogue, Captioning, Referring, REC, RES, GroundCap
GSVA Xia et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L ${}^{\blacklozenge}$ VQA, Segmentation, REC, RES
Lenna Wei et al. (2023) ${}^{\blacktriangle}$ CLIP ViT-L ${}^{\blacklozenge}$ VQA, Captioning, REC
LISA++ Yang et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L ${}^{\blacklozenge}$ Visual Dialogue, Captioning, RES
LLaVA-G Zhang et al. (2023d) ${}^{\blacklozenge}$ CLIP ViT-L OpenSeeD, ${}^{\blacklozenge}$ Visual Dialogue, REC, RES, Grounding
PixelLLM Xu et al. (2023a) ${}^{\blacktriangle}$ EVA ViT-L ${}^{\bigstar}$ Referring, REC, RES, GroundCap
PixelLM Ren et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L - Visual Dialogue, RES
VistaLLM Pramanick et al. (2023) ${}^{\blacklozenge}$ EVA - Visual Dialogue, VQA, Referring, REC, RES, GroundCap
ChatterBox Tian et al. (2024b) ${}^{\blacktriangle}$ CLIP ViT-L ${}^{\bigstar}$, ${}^{\blacklozenge}$ Visual Dialogue, Referring, REC, GroundCap
GELLA Qi et al. (2024) ${}^{\blacktriangle}$ CLIP ViT-L ${}^{\blacklozenge}$ Segmentation, RES, GroundCap
PaLI-3 Chen et al. (2023h) ${}^{\blacklozenge}$ SigLIP ViT-g ${}^{\blacklozenge}$ VQA, Captioning, Retrieval, RES
Table 2: Summary of MLLMs with components specifically designed for visual grounding and region-level understading. For each model, we indicate the LLM used in its best configuration, in some cases initialized with the weights of a pre-trained MLLM, and any supporting models used to perform the task ($\blacklozenge$: fine-tuning; $\blacktriangle$: fine-tuning with PEFT techniques; $\bigstar$: frozen). Gray color indicates models not publicly available.
@@ 3.1 Visual Grounding
Embedding-as-Region. Another solution is to read input regions through region encoders and provide the output regions as embeddings extracted from the last layer of the MLLM to a decoder. For input regions, GLaMM Rasheed et al. (2023), GPT4RoI Zhang et al. (2023f), ASM Wang et al. (2023d) and ChatterBox Tian et al. (2024b) leverage features of the image encoder to perform ROI align on the bounding box, whereas PVIT Chen et al. (2023a) exploits RegionCLIP Zhong et al. (2022). PixelLLM Xu et al. (2023a) and LLaVA-G Zhang et al. (2023d) use the prompt encoder of SAM Kirillov et al. (2023) and Semantic-SAM Li et al. (2023d) respectively. For output regions, LISA Lai et al. (2023), GLaMM, GSVA Xia et al. (2023b), NeXt-Chat Zhang et al. (2023a), and LISA++ Yang et al. (2023b) send the embedding corresponding to special tokens to the mask decoder of SAM, LLaVA-G to OpenSeeD Zhang et al. (2023c), Lenna Wei et al. (2023) to Grounding-DINO Liu et al. (2023i), and PixelLM Ren et al. (2023b) to a custom lightweight pixel decoder.
Differently, ContextDET Zang et al. (2023) introduces a decoder that receives the latent embedding of the noun with learnable queries, performs a cross-attention with image features, and then uses a segmentation head. ChatterBox Tian et al. (2024b) combines features from the iTPN-B encoder Tian et al. (2023) and the MLLM and provides them to the DINO detector Zhang et al. (2022a). GELLA Qi et al. (2024) presents a fusion module in Mask2Former Cheng et al. (2022) to propose masks based on multi-modal image features and an association module to assign latent embeddings to them. PaLI-3 Chen et al. (2023h) converts embeddings into segmentation masks through a VQ-VAE Van Den Oord et al. (2017) decoder.
@@
GILL Koh et al. (2023a) ${}^{\bigstar}$ CLIP ViT-L SD ${}^{\bigstar}$ Visual Dialogue, Retrieval, Image Generation
Emu Sun et al. (2023b) ${}^{\blacklozenge}$ EVA ViT-g SD ${}^{\blacklozenge}$ Visual Dialogue, VQA, Captioning, Image Generation
SEED Ge et al. (2023a) ${}^{\blacktriangle}$ EVA ViT-g SD ${}^{\bigstar}$ VQA, Captioning, Image Generation
DreamLLM Dong et al. (2023) ${}^{\blacklozenge}$ CLIP ViT-L SD ${}^{\bigstar}$ Visual Dialogue, VQA, Captioning, Image Generation, Interleaved Generation
LaVIT Jin et al. (2023) ${}^{\blacklozenge}$ EVA ViT-g SD ${}^{\blacklozenge}$ VQA, Captioning, Image Generation
MGIE Fu et al. (2024) ${}^{\bigstar}$ CLIP ViT-L SD ${}^{\blacklozenge}$ Image Editing
TextBind Li et al. (2023e) ${}^{\blacklozenge}$ EVA ViT-g SD ${}^{\bigstar}$ Visual Dialogue, VQA, Captioning, Image Generation
Kosmos-G Pan et al. (2023) ${}^{\lozenge}$ CLIP ViT-L SD ${}^{\bigstar}$ Image Generation, Compositional Image Generation
MiniGPT-5 Zheng et al. (2023) ${}^{\blacktriangle}$ EVA ViT-g SD ${}^{\bigstar}$ Visual Dialogue, Image Generation, Interleaved Generation
SEED-LLaMA Ge et al. (2023b) ${}^{\blacklozenge}$ EVA ViT-g SD ${}^{\bigstar}$ Visual Dialogue, VQA, Captioning, Image Generation, Interleaved Generation
CoDi-2 Tang et al. (2023) ${}^{\blacktriangle}$ ImageBind SD ${}^{\bigstar}$ Visual Dialogue, Audio Understanding, Image Generation, Image Editing
Emu2 Sun et al. (2023a) ${}^{\blacklozenge}$ EVA ViT-E SD ${}^{\blacklozenge}$ Visual Dialogue, VQA, Captioning, Image Generation, Image Editing
LLMGA Xia et al. (2023a) ${}^{\blacklozenge}$ CLIP ViT-L SD ${}^{\blacklozenge}$ Visual Dialogue, VQA, Image Generation, Image Editing
SmartEdit Huang et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L ${}^{\blacklozenge}$ Image Editing
VL-GPT Zhu et al. (2023b) ${}^{\blacktriangle}$ CLIP ViT-L SD ${}^{\bigstar}$ Visual Dialogue, VQA, Captioning, Image Generation, Image Editing
MM-Interleaved Tian et al. (2024a) ${}^{\blacklozenge}$ CLIP ViT-L SD ${}^{\blacklozenge}$ VQA, Captioning, REC, Image Generation, Interleaved Generation
JAM Aiello et al. (2024) ${}^{\blacklozenge}$ ${}^{\blacklozenge}$ Image Generation, Interleaved Generation

### v2

@@ 2.3 Vision-to-Language Adapters
Q-Former. It is a Transformer-based model proposed in BLIP-2 Li et al. (2023g) and then used in several other approaches Chen et al. (2023d); Dai et al. (2023); Hu et al. (2024). It is characterized by its adaptable architecture, which consists of two Transformer blocks that share mutual self-attention layers, facilitating the alignment process between visual and textual representations. It involves a set of learnable queries that interact within the self-attention layers and interface with visual features via a cross-attention mechanism. Textual and visual elements communicate via shared self-attention within the modules.
@@
Additional Cross-Attention Layers. This approach has been proposed in Flamingo Alayrac et al. (2022) with the integration of dense cross-attention blocks among the existing pre-trained layers of the LLM. The newly added layers are often combined with a zero-initialized tanh-gating mechanism to ensure that, upon initialization, the conditioned model acts as its original version. The use of additional cross-attention layers imposes the need to train them from scratch, increasing the number of trainable parameters compared to other alternatives. To reduce computational complexity, this strategy is usually paired with a Perceiver-based component Jaegle et al. (2021) that reduces the number of visual tokens before they are fed to the LLM. Since its introduction, several models Awadalla et al. (2023); Chen et al. (2023b); Laurençon et al. (2024); Li et al. (2023b) employ this technique to connect the visual modality with the underlying LLM, demonstrating enhanced training stability and improved performance.
@@ 2.4 Multimodal Training
Single-Stage Training. This possibility is explored by LLaMA-Adapter Gao et al. (2023) which introduces additional trainable parameters to encapsulate the visual knowledge and manage text-only instruction learning at the same time. To achieve this, the model undergoes joint training using image-text pairs and instructions, operating on separate parameters. Concurrently, the model proposed in Koh et al. (2023b) adapts the final loss function by incorporating two contrastive losses for image-text retrieval. During the training, only three linear layers are updated. On a different line, Kosmos-1 Huang et al. (2023b) considers a frozen visual backbone and trains the language model of 1.3B parameters from scratch.
Flamingo Alayrac et al. (2022) and its open source variants Awadalla et al. (2023); Laurençon et al. (2024), instead, train the cross-attention layers and the Perceiver-based component to connect the visual features with the frozen LLM blocks. Additionally, Otter Li et al. (2023b) extends Flamingo’s training to increment its in-context capabilities. Given the amount of training data currently available, approaches like SPHINX-X Gao et al. (2024) opt to perform a single all-in-one training stage in which to update all model components, possibly also using text-only data to preserve the conversational capabilities of the LLM.
@@
Training Data. During the first (or single) training stage, the datasets predominantly consist of large-scale, publicly available, and uncurated data. For instance, the Conceptual Captions 3M (CC3M) dataset Sharma et al. (2018) is composed of 3M images paired with textual captions specifically designed for image captioning systems. Unlike the widely-used and curated MS-COCO Lin et al. (2014) dataset, which serves similar purposes, images and captions in CC3M are gathered from the web, showcasing a broader spectrum of styles and content. Similarly, the LAION family Schuhmann et al. (2021, 2022) represents an extended collection of non-curated image-text pairs sourced from web pages, providing a rich resource for pre-training multimodal language models. Additionally, the COYO-700M Byeon et al. (2022) dataset stands out as a significant resource, containing 747M image-text pairs. Notably, each alt-text in COYO-700M is linked to an image within HTML documents. Furthermore, DataComp Gadre et al. (2023) presents an extensive pool of 12.8B filtered image-text pairs sourced from common crawl.
It is important to highlight the distinction between datasets used in the initial phase of training, which typically comprise large-scale, uncurated data, and those selected for refinement in subsequent stages. While the former emphasizes diversity and scale, the latter focuses on specificity and task relevance, facilitating a more tailored approach to model optimization. Especially in single-training stage approaches, certain methods Alayrac et al. (2022); Laurençon et al. (2024) also leverage interleaved datasets, which contain images interleaved with text coming from the web, aiming to augment the dataset size for large models Hoffmann et al. (2022). Images within these datasets can be positioned at the beginning or in the middle of a sentence, allowing models to support arbitrarily interleaved sequences of images and text as input, thereby enhancing flexibility in input formats by blending textual and visual elements. Among these datasets, the most used are WebLI Chen et al. (2023j), composed of 10B images and image-text pairs, and MMC4 Zhu et al. (2023d), an extension of the text-only C4 Raffel et al. (2020) dataset composed of 365M documents and 156B tokens relatives to different concepts, and OBELICS Laurençon et al. (2024), an open and curated collection of interleaved image-text web documents, containing 141M documents, 115B text tokens, and 353M images.
In the context of visual instruction tuning, which constitutes the second training stage for MLLMs, the available amount of data is limited. This limitation is mainly due to the creation process which is time-consuming and less well-defined. In this phase, different datasets are used to improve performances on a series of downstream tasks. Among them, LLaVA-Instruct Liu et al. (2023e) is a collection of GPT-4 generated multimodal instruction-following data. It comprises 158k unique language-image descriptions, spanning various types of tasks including 58k conversations, 23k detailed descriptions, and 77k complex reasoning. Similarly, LRV-Instruction Liu et al. (2023c) initially consisted of 400k visual instructions generated by GPT-4, and more recently, it has been updated with an additional set of 300k visual instructions. To enhance robustness in instruction tuning, LRV-Instruct also includes negative instructions organized across three semantic levels, showing that instruct-tuned MLLMs on this dataset suffer less from hallucination compared to the original versions. Moreover, LLaVAR Zhang et al. (2023i) considers publicly available OCR tools to collect results on 422k text-rich images from the LAION dataset. The pipeline first collects 422k noisy text-rich images and then extracts the text through OCR models. With the help of GPT-4, the results and captions are used to create 16k conversations, also including specific questions to create complex instructions which can be helpful to boost performance on new tasks.
@@
ContextDET Zang et al. (2023) OPT-6.7B Swin-B - Visual Dialogue, VQA, Captioning, Detection, REC, RES
DetGPT Pi et al. (2023) Vicuna-13B EVA ViT-g G-DINO Visual Dialogue, Detection
VisionLLM Wang et al. (2023e) Alpaca-7B Intern-H Deformable-DETR VQA, Captioning, Detection, Segmentation, REC
BuboGPT Zhao et al. (2023c) Vicuna-7B EVA ViT-g RAM, G-DINO, SAM Visual Dialogue, Audio Understanding, Captioning, GroundCap
ChatSpot Zhao et al. (2023b) Vicuna-7B CLIP ViT-L - Visual Dialogue, VQA, Captioning, Referring
GPT4RoI Zhang et al. (2023g) LLaVA-7B OpenCLIP ViT-H - Visual Dialogue, VQA, Captioning, Referring
ASM Wang et al. (2023d) Husky-7B EVA ViT-g - VQA, Captioning, Referring
LISA Lai et al. (2023) LLaVA-13B CLIP ViT-L SAM Visual Dialogue, Captioning, RES
PVIT Chen et al. (2023a) LLaVA-7B CLIP ViT-L RegionCLIP Visual Dialogue, VQA, Captioning, Referring
GLaMM Rasheed et al. (2023) Vicuna-7B OpenCLIP ViT-H SAM Visual Dialogue, Captioning, Referring, REC, RES, GroundCap
Griffon Zhan et al. (2023) LLaVA-13B CLIP ViT-L - REC, Detection, Phrase Grounding
LLaFS Zhu et al. (2023c) CodeLLaMA-7B CLIP RN50 - Few-Shot Segmentation
NExT-Chat Zhang et al. (2023a) Vicuna-7B CLIP ViT-L SAM Visual Dialogue, Captioning, Referring, REC, RES, GroundCap
GSVA Xia et al. (2023b) LLaVA-13B CLIP ViT-L SAM VQA, Segmentation, REC, RES
Lenna Wei et al. (2023) LLaVA-7B CLIP ViT-L G-DINO VQA, Captioning, REC
LISA++ Yang et al. (2023b) LLaVA-13B CLIP ViT-L SAM Visual Dialogue, Captioning, RES
LLaVA-G Zhang et al. (2023d) Vicuna-13B CLIP ViT-L OpenSeeD, S-SAM Visual Dialogue, REC, RES, Grounding
PixelLLM Xu et al. (2023a) FlanT5-XL-3B EVA ViT-L SAM Referring, REC, RES, GroundCap
PixelLM Ren et al. (2023b) LLaVA-7B CLIP ViT-L - Visual Dialogue, RES
VistaLLM Pramanick et al. (2023) Vicuna-13B EVA - Visual Dialogue, VQA, Referring, REC, RES, GroundCap
ChatterBox Tian et al. (2024b) LLaVA-13B CLIP ViT-L iTPN-B, DINO Visual Dialogue, Referring, REC, GroundCap
GELLA Qi et al. (2024) LLaVA-13B CLIP ViT-L Mask2Former Segmentation, RES, GroundCap
PaLI-3 Chen et al. (2023i) UL2-3B SigLIP ViT-g VQ-VAE VQA, Captioning, Retrieval, RES
Table 2: Summary of MLLMs with components specifically designed for visual grounding and region-level understanding. For each model, we indicate the LLM used in its best configuration, in some cases initialized with the weights of a pre-trained MLLM, and any supporting models used to perform the task ($\blacklozenge$: fine-tuning; $\blacktriangle$: fine-tuning with PEFT techniques; $\bigstar$: frozen). Gray color indicates models not publicly available.
@@ 3.1 Visual Grounding
Embedding-as-Region. Another solution is to read input regions through region encoders and provide the output regions as embeddings extracted from the last layer of the MLLM to a decoder. For input regions, GLaMM Rasheed et al. (2023), GPT4RoI Zhang et al. (2023g), ASM Wang et al. (2023d) and ChatterBox Tian et al. (2024b) leverage features of the image encoder to perform ROI align on the bounding box, whereas PVIT Chen et al. (2023a) exploits RegionCLIP Zhong et al. (2022). PixelLLM Xu et al. (2023a) and LLaVA-G Zhang et al. (2023d) use the prompt encoder of SAM Kirillov et al. (2023) and Semantic-SAM Li et al. (2023e) respectively. For output regions, LISA Lai et al. (2023), GLaMM, GSVA Xia et al. (2023b), NeXt-Chat Zhang et al. (2023a), and LISA++ Yang et al. (2023b) send the embedding corresponding to special tokens to the mask decoder of SAM, LLaVA-G to OpenSeeD Zhang et al. (2023c), Lenna Wei et al. (2023) to Grounding-DINO Liu et al. (2023i), and PixelLM Ren et al. (2023b) to a custom lightweight pixel decoder.
Differently, ContextDET Zang et al. (2023) introduces a decoder that receives the latent embedding of the noun with learnable queries, performs a cross-attention with image features, and then uses a segmentation head. ChatterBox Tian et al. (2024b) combines features from the iTPN-B encoder Tian et al. (2023) and the MLLM and provides them to the DINO detector Zhang et al. (2022a). GELLA Qi et al. (2024) presents a fusion module in Mask2Former Cheng et al. (2022) to propose masks based on multi-modal image features and an association module to assign latent embeddings to them. PaLI-3 Chen et al. (2023i) converts embeddings into segmentation masks through a VQ-VAE Van Den Oord et al. (2017) decoder.
@@
GILL Koh et al. (2023a) OPT-6.7B CLIP ViT-L SD v1.5 Visual Dialogue, Retrieval, Image Generation
Emu Sun et al. (2023b) LLaMA-13B EVA ViT-g SD v1.5 Visual Dialogue, VQA, Captioning, Image Generation
SEED Ge et al. (2023a) OPT-2.7B EVA ViT-g SD v1.4 VQA, Captioning, Image Generation
DreamLLM Dong et al. (2023) Vicuna-7B CLIP ViT-L SD v2.1 Visual Dialogue, VQA, Captioning, Image Generation, Interleaved Generation
LaVIT Jin et al. (2023) LLaMA-7B EVA ViT-g SD v1.5 VQA, Captioning, Image Generation
MGIE Fu et al. (2024) LLaVA-7B CLIP ViT-L SD v1.5 Image Editing
TextBind Li et al. (2023f) LLaMA-2-7B EVA ViT-g SD XL Visual Dialogue, VQA, Captioning, Image Generation
Kosmos-G Pan et al. (2023) Magneto-1.3B CLIP ViT-L SD v1.5 Image Generation, Compositional Image Generation
MiniGPT-5 Zheng et al. (2023) Vicuna-7B EVA ViT-g SD v2.1 Visual Dialogue, Image Generation, Interleaved Generation
SEED-LLaMA Ge et al. (2023b) LLaMA-2-13B EVA ViT-g SD unCLIP Visual Dialogue, VQA, Captioning, Image Generation, Interleaved Generation
CoDi-2 Tang et al. (2023) LLaMA-2-7B ImageBind SD unCLIP Visual Dialogue, Audio Understanding, Image Generation, Image Editing
Emu2 Sun et al. (2023a) LLaMA-33B EVA ViT-E SD XL Visual Dialogue, VQA, Captioning, Image Generation, Image Editing
LLMGA Xia et al. (2023a) LLaVA-13B CLIP ViT-L SD XL Visual Dialogue, VQA, Image Generation, Image Editing
SmartEdit Huang et al. (2023c) LLaVA-13B CLIP ViT-L SD Image Editing
VL-GPT Zhu et al. (2023b) LLaMA-7B CLIP ViT-L SD v1.5 Visual Dialogue, VQA, Captioning, Image Generation, Image Editing
MM-Interleaved Tian et al. (2024a) Vicuna-13B CLIP ViT-L SD v2.1 VQA, Captioning, REC, Image Generation, Interleaved Generation
JAM Aiello et al. (2024) LLaMA-7B - CM3Leon Image Generation, Interleaved Generation